# Deep-Graph-Sprints: Accelerated Representation Learning in Continuous-Time Dynamic Graphs

### v1

@@ 5.3 Low-latency methods
In this paper, we are interested in low latency methods. APAN (Wang et al., 2021) proposes to reduce the latency at inference time by decoupling the more costly graph operations from the inference module, and perform the more expensive $k$-hop message passing asynchronously. While APAN improves the latency at inference time, it potentially uses outdated information at inference time due to asynchronous update. This algorithm addresses the need to generate low-latency embeddings, similar to our Deep-Graph-Sprints methods. However, unlike our methods, APAN uses outdated information which could negatively affect its overall performance.
Moreover, towards reducing computational costs of GNNs, HashGNN (Wu et al., 2021) leverages MinHash to generate node embeddings suitable for the link prediction task, where nodes that results in the same hashed embedding are considered similar. SGSketch (Yang et al., 2022) is a streaming node embedding framework uses a mechanism to gradually forget outdated edges, achieving significant speedups. Differently than our approach SGSketch uses the gradual forgetting strategy to update the adjacency matrix and therefore only considers the graph structure.
Liu et al. (2019) propose an algorithm for graph streams that performs node representation updates in real-time by leveraging the embeddings of the 1-hop neighbors of a node of interest, and ignoring its attributes.
Chu & Lin (2024) propose ETSMLP, a model that leverages an exponential smoothing technique to model long-term dependencies in sequence learning. The model employs two learnable damped factors, to modify the influence of a smoothing factor and the current input, respectively. These factors enable the model to adjust the impact of past and current data adaptively. It treats the learnable and smoothing factors as complex numbers for a richer representation.
@@ 6 Conclusions
CTDGs are essential for representing connected and evolving systems. While methods to learn these graphs show significant potential for diverse applications, the computational and memory demands of existing approaches limit their feasibility in low-latency scenarios.
@@ Appendix A Appendix
Table 6 provide a detailed comparison of the number of learnable parameters between DGS and TGN-attn in the link prediction task.

### v3

@@ 5.2 Low-latency Graph Representation Learning
This section reviews methods for low-latency graph representation learning. For example, APAN (Wang et al., 2021) aims to reduce inference latency by decoupling expensive graph operations from the inference module, executing the costly $k$-hop message passing asynchronously. While APAN enhances inference efficiency, it may use outdated information due to its asynchronous updates, which could impact overall performance. In contrast, our method, Deep-Graph-Sprints, addresses latency without compromising the freshness of the information used.
Furthermore, Liu et al. (2019) present a real-time algorithm for graph streams that updates node representations based on the embeddings of 1-hop neighbors of a node of interest, and ignoring its attributes. Chu & Lin (2024) propose ETSMLP, a model that leverages an exponential smoothing technique to model long-term dependencies in sequence learning. However, ETSMLP is tailored specifically for sequence modeling and is not applicable to graph-based tasks out-of-the-box. Graph-Sprints (Eddin et al., 2023) offers a feature engineering approach that approximates random walks for low-latency graph feature extraction. Unlike our approach, Graph-Sprints requires extensive hand-crafting of features.
In addition to inference optimization, several methods address the reduction of computational costs in GNNs. HashGNN(Wu et al., 2021) employs MinHash to generate node embeddings suitable for link prediction tasks, grouping similar nodes based on their hashed embeddings. Another approach, SGSketch(Yang et al., 2022), introduces a streaming node embedding framework that gradually forgets outdated edges, leading to speed improvements. Unlike our approach, SGSketch primarily updates the adjacency matrix and focuses on the graph structure rather than incorporating additional node or edge attributes.
@@ 6 Conclusions
CTDGs are essential for representing connected and evolving systems. The computational and memory demands associated with performing lookups to sample multiple neighbors limit their feasibility in low-latency scenarios.
@@
Future work includes exploring alternative normalization functions or activation functions beyond softmax and incorporating advanced optimization algorithms such as Adam to replace the current use of stochastic gradient descent for updating DGS parameters during forward-mode AD. Additionally, a significant enhancement under consideration is enabling input-dependent adaptability for the parameters $\vec{\alpha}$ and $\vec{\beta}$, aiming to improve the model’s responsiveness to varying input features and enhance overall performance, similar to approaches used in gated recurrent neural networks.
@@ Appendix A Appendix
A.1 Comparing Graph-Sprints and Deep-Graph-Sprints
Deep-Graph-Sprints builds upon and addresses the limitations of Graph-Sprints, resulting in both notable similarities and key distinctions, as outlined in Table 4.
Table 4: Comparison of Graph-Sprints and Deep-Graph-Sprints
Aspect Graph-Sprints Deep-Graph-Sprints
Components Embedding + classifier Embedding + classifier
Input embeddings Hard-coded (feature engineered) Learnable
Optimization of embedding and classifier parameters Separate End-to-end
Embedding size Fixed and large (bucketing of raw features) Tuneable (learnable embeddings)
Embedding expressivity Scalar forgetting coefficients Vector forgetting coefficients
A.2 Data: Specifications and Characteristics
Table 5 outlines the specifications and key characteristics of the five datasets utilized in the experiments described in Section 4.
Wikipedia Mooc Reddit FI-A FI-B
#Nodes 9,227 7,047 10,984 $\approx$400,000 $\approx$10,000
#Edges 157,474 411,749 672,447 $\approx$500,000 $\approx$2,000,000
Positive labels 0.14% 0.98% 0.05% 2-5% 20-40%
Duration 30 days 29 days 30 days $\approx$300 days $\approx$600 days
Used split (%) 70-15-15 60-20-20 70-15-15 60-10-30 60-10-30
A.3 Hyperparameters Ranges
@@
A.4 Ablation Study
We evaluate the performance of different variants of the DGS method on the node classification task. We compare three distinct approaches:
  • DGS-bp: DGS with the typical truncated backpropagation strategy as used in SOTA methods such as TGN and JODIE.
  • DGS-sum: DGS using a divide-by-sum normalization technique, which simplifies the normalization process by summing activations across the network. Equation 3 illustrates the process for a single partition in the $W$ matrix (analogous to a single softmax operation in our proposed approach).
    $\vec{E}_{i}=\frac{W_{i}\vec{F_{t}}}{\sum(W_{i}\vec{F_{t}})+\varepsilon}$ (3)
  • DGS (proposed): Our proposed method employs forward-mode AD and utilizes softmax normalization.
As shown in Table 7, our proposed DGS method outperforms both DGS-bp and DGS-sum, demonstrating the advantage of using forward-mode AD and softmax normalization for this task.
Table 7: Node classification performance comparison across DGS variants (DGS-bp, DGS-sum, and the proposed DGS method). Detailed descriptions of these variants are provided in Appendix A.4.
Method AUC ± std
Wikipedia Mooc Reddit
DGS-sum 84.5 ± 9.1 71.3 ± 3.9 53.7 ± 7.4
DGS-bp 85.8 ± 0.4 72.6 ± 13.2 63.2 ± 0.7
DGS (proposed) 89.2 ± 2.2 78.7 ± 0.6 68.0 ± 1.9
A.5 Comparison of Inference Speed
We evaluate and compare the inference speed of two different methods: DGS (proposed) and DGS-sum. The comparison is based on their performance across various datasets. Table 8 summarizes the average inference times, in seconds, along with their standard deviations. The results indicate that both methods exhibit similar inference speeds across the different datasets.
Table 8: Comparison of inference speed using different normalization functions
Data DGS (proposed) DGS-sum
Wikipedia 0.24 $\pm$ 0.002 0.25 $\pm$ 0.009
Reddit 0.24 $\pm$ 0.003 0.23 $\pm$ 0.0004
Mooc 0.28 $\pm$ 0.003 0.27 $\pm$ 0.0003
A.6 Comparison of Learnable Parameter Counts
Table 9 provides a comprehensive comparison of the learnable parameter counts between DGS and TGN-attn for the link prediction task across inductive and transductive settings, underscoring differences in model complexity.